# AgentLeak: A Full-Stack Benchmark for Privacy Leakage in Multi-Agent LLM Systems

### v1

@@ VII-E Future Research Directions
Privacy-Aware Coordination Protocols. Selective disclosure at the framework level—agents share minimum necessary context while protecting sensitive fields—is the critical near-term priority, with open questions around formally verifiable data-flow policies and dynamic workflow management.
Differential Privacy and MPC. Differential privacy on inter-agent messages and secure multi-party computation could enable coordination without exposing plaintext, though budget selection, composition, and performance overhead remain open challenges.
@@
Adversarial Robustness. The 80% attack success rate for coordination attacks (F4) motivates defenses specifically targeting inter-agent trust exploitation—absent in single-agent settings.
@@ VIII Conclusion
We presented AgentLeak, a benchmark for evaluating privacy leakage across multi-agent LLM systems. The results convey a clear pattern: through 4,979 execution traces across five production LLMs (GPT-4o, GPT-4o-mini, Claude-3.5-Sonnet, Mistral Large, Llama 3.3 70B), we found that internal channels—the messages agents send each other, the memory they share—account for most privacy violations.
@@
Second, Claude-3.5-Sonnet, which emphasizes safety alignment in its design, achieves the lowest rates on both external (3.3% on C1) and internal (28.1% on C2) channels, suggesting that model-level safety training may transfer to internal channel protection. However, we cannot isolate the contribution of safety alignment from other architectural differences between models (e.g., training data, RLHF methodology, model size). The correlation is suggestive but not causal.
Third, output-only audits miss 41.7% of privacy violations. Organizations deploying agentic systems in regulated domains should audit all channels, not just final outputs.
The privacy gap we identify is structural but not fundamental. Our prototype interceptor reduces internal leakage from 31.5% to 2.4% on a 120-scenario subset, with a 4.7-point TSR reduction, illustrating the privacy–utility tradeoff. Addressing the gap requires architectural changes at the framework level: message-level sanitization intercepting inter-agent communication, memory access controls with field-level permissions, and full-channel auditing as a default capability.
We encourage the research community to extend the benchmark with additional scenarios, frameworks, and defense mechanisms as the multi-agent ecosystem continues to evolve.
@@ Appendix A Reproducibility Details
Acknowledgments
AI-assisted editing disclosure: We used ChatGPT (OpenAI, GPT-4) and Claude (Anthropic, Claude 3.5 Sonnet) to assist with grammar and style edits across the manuscript (Abstract, Introduction, Related Work, Problem Definition, Benchmark Design, Attack Taxonomy, Evaluation, Discussion, and Conclusion). These systems were used only for language polishing (clarity, grammar, and phrasing) and did not generate new technical content, results, or claims.

### v2

@@ VII-E Future Research Directions
Privacy-Aware Coordination Protocols. Selective disclosure at the framework level (agents share minimum necessary context while protecting sensitive fields) is the critical near-term priority, with open questions around formally verifiable data-flow policies and dynamic workflow management.
Differential Privacy and Secure Multi-Party Computation (MPC). Differential privacy on inter-agent messages and secure multi-party computation could enable coordination without exposing plaintext [38, 39, 40], though budget selection, composition, and performance overhead remain open challenges.
@@
Adversarial Robustness. The 80% attack success rate for coordination attacks (F4) motivates defenses specifically targeting inter-agent trust exploitation, absent in single-agent settings.
@@ VIII Conclusion
The gap, however, is not irreparable. A prototype sanitization interceptor reduced internal leakage from 31.5% to 2.4% at the cost of a 4.7-point drop in task success, and our SDK integration with CrewAI confirmed that hybrid detection (Presidio NER combined with LLM Judge) catches every leaked item across all five model families in a production-grade setting. Importantly, these mitigations operate at the framework layer rather than requiring model retraining, suggesting that privacy-preserving coordination is achievable without sacrificing model capability. Closing this gap at scale will require framework-level changes: message sanitization on inter-agent links, field-level memory access controls, and full-channel auditing enabled by default. The research directions outlined in Section VII, from privacy-aware coordination protocols to cryptographic safeguards and broader topology coverage, chart a concrete path forward. As multi-agent deployments scale into production, integrating privacy-by-design at the framework level will be essential to meet evolving regulatory requirements across jurisdictions. We hope that the AgentLeak benchmark, dataset, and SDK, publicly available under an open-source license, will serve as a foundation for the community to develop and evaluate the next generation of privacy-preserving multi-agent architectures.
@@ Appendix A Reproducibility Details
Acknowledgment
AI-assisted editing disclosure: The authors used ChatGPT (OpenAI, GPT-4) and Claude 3.5 Sonnet (Anthropic) to assist with grammar and style edits across the manuscript (Abstract, Introduction, Related Work, Problem Definition, Benchmark Design, Attack Taxonomy, Evaluation, Discussion, and Conclusion). These systems were used only for language polishing (clarity, grammar, and phrasing) and did not generate new technical content, results, or claims.